2x+y=6
5x+4y=0
im not sure how to solve these questions among others

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

2x+y=6
5x+4y=0
-----------
There's more than one method.
You can use substitution:
2x+y=6
--> y = 6-2x
Sub for y in the 2nd eqn
5x + 4(6-2x) = 0
Solve for x.
Then sub for x into either eqn and find y.
---------------
Elimination works, too:
2x+y=6
times 4 --> 8x+4y = 24
-----------
8x+4y = 24
5x+4y = 0
----------- Subtract
3x = 24
x = 8
Then sub and find y.

Answer by ewatrrr(24785)

 
Hi,
Solving using the Substitution Method
2x+y=6 OR y = -2x+6
5x+4y=0
5x + 4(-2x+6) = 0
-3x = -24
x = 8 and y = -10
Ordered pair (8,-10) is the solution for this system of EQs
CHECKING our Answer***Using Elimination Method
2x+y=6
5x+4y=0
-8x-4y=-24 |Multiplying 1st EQ thru by -4, so as to eliminate y
5x +4y = 0
-3x = -24
x = 8 and y = -10 (y = -2x+6)